OpenStudy (anonymous):

Simplifying Algebraic Expressions... 9x^2-14x^2+20x^2

OpenStudy (xxdragoonxx):

9x^2-14x^2+20x^2 = 9x^2+-14x^2=-5x^2-5x^2+20x^2 = -5x^2+20x^2=15x^2